if sinx+sin2x=1 then cos2x +cos4x=?

Grade:

4 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

sinx=1-sin2x=cos2x

Use this to solve...

Shubham
9 years ago
sinx=1-sin2x=cos2x
therefore.. sin2x=cos4x
cos2x+cos4x=1
 
Shubham
9 years ago
sinx=1-sin2x=cos2x
therefore.. sin2x=cos4x
cos2x+cos4x=1
 
Shubham
9 years ago
sinx=1-sin2x=cos2x
therefore.. sin2x=cos4x
cos2x+cos4x=1